# Distributional Reduction: Unifying Dimensionality Reduction and Clustering with Gromov-Wasserstein

### v2

@@ A.3 Necessary and sufficient condition
The latter problem is a co-optimal transport problem [60], and, since it is a bilinear problem, there is an optimal solution $({\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)})$ such that both ${\mathbf{T}}^{(1)}$ and ${\mathbf{T}}^{(2)}$ are in an extremal point of $\operatorname{DS}$ which is the space of $N\times N$ permutation matrices by Birkhoff’s theorem [6]. This point was already noted by [37] but we recall the proof for completeness. We note $L_{ijkl}=L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})$ and consider an optimal solution $({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ of $\min_{{\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}\in\operatorname{DS}}\phi({\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}):=\sum_{ijkl}L_{ijkl}T^{(1)}_{ik}T^{(2)}_{jl}$. Consider the linear problem $\min_{{\mathbf{T}}\in\operatorname{DS}}\phi({\mathbf{T}},{\mathbf{T}}_{\star}^{(2)})$. Since it is a linear over the space of doubly stochastic matrices it admits a permutation matrix ${\mathbf{P}}^{(1)}$ as optimal solution. Also $\phi({\mathbf{P}}^{(1)},{\mathbf{T}}_{\star}^{(2)})\leq\phi({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ by optimality. Now consider the linear problem $\min_{{\mathbf{T}}\in\operatorname{DS}}\phi({\mathbf{P}}^{(1)},{\mathbf{T}})$, in the same it admits a permutation matrix ${\mathbf{P}}^{(2)}$ as optimal solution and by optimality $\phi({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})\leq\phi({\mathbf{P}}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ thus $\phi({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})\leq\phi({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ which implies that $({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})$ is an optimal solution. Combining with eq. 27 we get
@@
where $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R},\ h:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $g:\mathbb{R}^{N\times d}\to\mathbb{R}$ which is permutation invariant [9], then $\mathcal{C}$ is stable under permutation. Indeed, permuting the rows and columns of ${\mathbf{C}}_{Z}({\mathbf{Z}})$ by $\sigma$ is equivalent to considering the similarity ${\mathbf{C}}_{Z}({\mathbf{Y}})$, where ${\mathbf{Y}}=(\bm{z}_{\sigma(1)},\cdots,\bm{z}_{\sigma(n)})^{\top}$. Moreover, most similarities in the target space considered in DR take the form eq. 32: $\langle\Phi(\bm{z}_{i}),\Phi(\bm{z}_{j})\rangle_{\mathcal{H}}$ (kernels such as in spectral methods with $\Phi=\operatorname{id}$), $f(\bm{z}_{i},\bm{z}_{j})/\sum_{nm}f(\bm{z}_{n},\bm{z}_{m})$ (normalized similarities such as in SNE and t-SNE). Also note that the condition on $\mathcal{C}=\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$ of proposition A.5 is met as soon as ${\mathbf{C}}_{Z}:\mathbb{R}^{N\times d}\to\mathbb{R}^{N\times N}$ is permutation equivariant.
Appendix B Generalized Semi-relaxed Gromov-Wasserstein is a divergence
Remark B.1 (Weak isomorphism).
According to the notion of weak isomorphism in [17], for a graph $({\mathbf{C}},{\mathbf{h}})$ with corresponding discrete measure $\mu=\sum_{i}h_{i}\delta_{x_{i}}$, two nodes $x_{i}$ and $x_{j}$ are the “same” if they have the same internal perception i.e $C_{ii}=C_{jj}=C_{ij}=C_{ji}$ and external perception $\forall k\neq(i,j),C_{ik}=C_{jk},C_{ki}=C_{kj}$. So two graphs $({\mathbf{C}}_{1},{\mathbf{h}}_{1})$ and $({\mathbf{C}}_{2},{\mathbf{h}}_{2})$ are said to be weakly isomorphic, if there exist a canonical representation $({\mathbf{C}}_{c},{\mathbf{h}}_{c})$ such that $\operatorname{card}(\operatorname{supp}({\mathbf{h}}_{c}))=p\leq n,m$ and $M_{1}\in\left\{0,1\right\}^{n\times p}$ (resp. $M_{2}\in\left\{0,1\right\}^{m\times p}$) such that for $k\in\left\{1,2\right\}$
@@ Appendix B Generalized Semi-relaxed Gromov-Wasserstein is a divergence
We first emphasize a simple result.
Proposition B.2.
@@
()($\Rightarrow$)( ⇒ ) As $\operatorname{srGW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$ there exists ${\mathbf{T}}\in\mathcal{U}({\mathbf{h}},\overline{{\mathbf{h}}})$ such that $E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=0$ hence $\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$. Using proposition B.2, $\operatorname{GW}_{L_{2}}=0$ hence using Theorem 18 in [17], it implies that $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ are weakly isomorphic.
@@
B.1 About trivial solutions of semi-relaxed GW when $L$ is not a proper divergence
@@
Appendix C Clustering properties: Proof of theorem 4.1
@@
We recall that $F({\mathbf{T}},\overline{{\mathbf{C}}}):=E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ and $G({\mathbf{T}})=F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))$. From lemma C.1 we know that $\overline{{\mathbf{C}}}({\mathbf{T}})$ is a minimizer of $\overline{{\mathbf{C}}}\to F({\mathbf{T}},\overline{{\mathbf{C}}})$ hence it satisfies the first order conditions $\partial_{2}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))=0$. Every quantity is differentiable on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$. Hence, taking the derivative of $G$ and using the first order conditions
@@
We will found the expression of this gradient. In the proof of lemma C.1 we have seen that
@@
From lemma C.1 we have the expression (since ${\mathbf{P}}\in{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$)
@@
Since ${\mathbf{U}}$ is symmetric we also have ${\mathbf{U}}\bm{1}_{N}={\mathbf{h}}_{X}$ and similarly we have the same result for ${\mathbf{V}}$. Overall ${\mathbf{V}}^{\top}\bm{1}_{N}={\mathbf{U}}^{\top}\bm{1}_{N}$ and ${\mathbf{V}}\bm{1}_{N}={\mathbf{U}}\bm{1}_{N}$. Since ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND we can apply lemma C.3 below which proves that $-\operatorname{Tr}({\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))-\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}))+2\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))\leq 0$ and consequently that $G$ is concave on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$. We now use the continuity of $G$ to prove that it is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$.
@@
and hence $G$ is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. This proves theorem 4.1. Indeed the minimization of ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})\to\min_{\overline{{\mathbf{C}}}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ is a minimization of a concave function over a polytope, hence admits an extremity of ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$ as minimizer. But these extremities are membership matrices as they can be described as $\{\operatorname{diag}({\mathbf{h}}_{X}){\mathbf{P}}:{\mathbf{P}}\in\{0,1\}^{N\times n},{\mathbf{P}}^{\top}\bm{1}_{n}=\bm{1}_{N}\}$ [11]. ∎
Lemma C.3.
@@
Appendix D Algorithmic details
@@
D.1 Objective function and gradient computation.
@@
Let consider any matrices ${\mathbf{C}}\in\mathbb{R}^{n\times n}$, $\overline{{\mathbf{C}}}\in\mathbb{R}^{m\times m}$, and a probability vector ${\mathbf{h}}\in\Sigma_{n}$. In all our use cases, we considered inner losses $L:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ which can be decomposed following Proposition 1 in [58]. Namely we assume the existence of functions $f_{1},f_{2},h_{1},h_{2}$ such that
@@
$\begin{split}L_{2}(a,b)=(a-b)^{2}&\implies f_{1}(a)=a^{2},\quad f_{2}(b)=b^{2},\quad h_{1}(a)=a,\quad h_{2}(b)=2b\\ L_{KL}(a,b)=a\log\frac{a}{b}-a+b&\implies f_{1}(a)=a\log a-a,\quad f_{2}(b)=b,\quad h_{1}(a)=a,\quad h_{2}(b)=\log b\\ \end{split}$ (L2)
@@
Inspired from the work of [65], we propose implementations of $\operatorname{srGW}$ that can leverage the low-rank nature of ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$. Let us assume that both structures can be exactly decomposed as follows, ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{A}}_{1}{\mathbf{A}}_{2}^{\top}$ where ${\mathbf{A}}_{1},{\mathbf{A}}_{2}\in\mathbb{R}^{N\times r}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top}$ with ${\mathbf{B}}_{1},{\mathbf{B}}_{2}\in\mathbb{R}^{n\times s}$, such that $r<<N$ and $s<<n$, can differ respectively from respective dimensions $p$ and $d$ (e.g. for used squared Euclidean distance matrices $r=p+2$ and $s=d+2$ ). For both inner losses $L$ we make use of the following factorization:
$L=L_{2}$: Computing the first term $F_{1}$ coming for the optimized second marginal can benefit from being factored if $d^{2}<<n$. Indeed, as $f_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))={\mathbf{C}}_{Z}({\mathbf{Z}})^{2}=({\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top})\odot({\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top})$, one can use the flattened out product operator described in [65, Section 5], to compute ${\mathbf{C}}_{Z}({\mathbf{Z}})^{2}{\mathbf{T}}^{\top}\bm{1}_{N}={\mathbf{x}}$ in $O(min(n^{2},ns^{2}))$. This way $F_{1}({\mathbf{T}})$ results from stacking $N$ times ${\mathbf{x}}$ in $O(1)$ operations for a total number of computions of $N+O(min(n^{2},ns^{2}))$.And its scalar product with ${\mathbf{T}}$ to compute the loss comes down to $O(Nn)$ additional operations.
Then computing $F_{2}({\mathbf{T}})$ and its scalar product with ${\mathbf{T}}$ can be done following the development of [65, Section 3] for the corresponding GW problem, in $O(Nn(r+s)+rs(N+n))$ operations. So the overall complexity at is $O(Nn(r+s)+rs(N+n)+min(n^{2},ns^{2}))$.
$L=L_{KL}$: In this setting $f_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))={\mathbf{C}}_{Z}({\mathbf{Z}})$ and $h_{1}({\mathbf{C}}_{X}({\mathbf{X}}))={\mathbf{C}}_{X}({\mathbf{X}})$ naturally preserves the low-rank nature of input matrices, but $h_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))=\log({\mathbf{C}}_{Z}({\mathbf{Z}}))$ does not. So computing the first term $F_{1}$, can be performed following this paranthesis order $\bm{1}_{N}((\bm{1}_{N}^{\top}{\mathbf{T}}){\mathbf{A}}_{1})){\mathbf{A}}_{2}^{\top})$ in $O(N(n+s))$ operations. While the second term $F_{2}$ should be computed following this order ${\mathbf{A}}_{1}(({\mathbf{A}}_{2}^{\top}{\mathbf{T}})\log({\mathbf{C}}_{Z}({\mathbf{Z}})))$ in $O(Nnr+rn^{2})$ operations. While their respective scalar product can be computed in $O(Nn)$. So the overall complexity is $O(Nnr+n^{2}r)$.
@@
D.2 Solvers.
We develop next our extension of both the Mirror Descent and Conditional Gradient solvers first introduced in [81], for any inner loss $L$ that decomposes as in equation 74 .
@@
where ${\mathbf{K}}^{(i)}=\exp\left(\ \nabla_{{\mathbf{T}}}F({\mathbf{T}}^{(i)})-\varepsilon\log({\mathbf{T}}^{(i)})\right)$ and $\varepsilon>0$ is an hyperparameter to tune. Proposition 3 and Lemma 7 in [80, Chapter 6] provides that the Mirror-Descent algorithm converges to a stationary point non-asymptotically when $L=L_{2}$. A quick inspection of the proof suffices to see that this convergence holds for any losses $L$ satisfying equation 74, up to adaptation of constants involved in the Lemma.
@@
This algorithm, known to converge to local optimum [39], iterates over the 3 steps summarized in Algorithm 1: .
@@
Observe that this objective function can be developed as a second order polynom $g(\gamma)=a\gamma^{2}+b\gamma+c$. To find an optimal $\gamma$ it suffices to express coefficients $a$ and $b$ to conclude using Algorithm 2 in [78].
@@
Appendix E Appendix of experimental section

### v3

@@ B.3 Necessary and sufficient condition
The latter problem is a co-optimal transport problem (Redko et al., 2020), and, since it is a bilinear problem, there is an optimal solution $({\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)})$ such that both ${\mathbf{T}}^{(1)}$ and ${\mathbf{T}}^{(2)}$ are in an extremal point of $\operatorname{DS}$ which is the space of $N\times N$ permutation matrices by Birkhoff’s theorem (Birkhoff, 1946). This point was already noted by Konno (1976) but we recall the proof for completeness. We note $L_{ijkl}=L([{\mathbf{C}}_{1}]_{ij},[{\mathbf{C}}_{2}]_{kl})$ and consider an optimal solution $({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ of $\min_{{\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}\in\operatorname{DS}}\phi({\mathbf{T}}^{(1)},{\mathbf{T}}^{(2)}):=\sum_{ijkl}L_{ijkl}T^{(1)}_{ik}T^{(2)}_{jl}$. Consider the linear problem $\min_{{\mathbf{T}}\in\operatorname{DS}}\phi({\mathbf{T}},{\mathbf{T}}_{\star}^{(2)})$. Since it is a linear over the space of doubly stochastic matrices it admits a permutation matrix ${\mathbf{P}}^{(1)}$ as optimal solution. Also $\phi({\mathbf{P}}^{(1)},{\mathbf{T}}_{\star}^{(2)})\leq\phi({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ by optimality. Now consider the linear problem $\min_{{\mathbf{T}}\in\operatorname{DS}}\phi({\mathbf{P}}^{(1)},{\mathbf{T}})$, in the same it admits a permutation matrix ${\mathbf{P}}^{(2)}$ as optimal solution and by optimality $\phi({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})\leq\phi({\mathbf{P}}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ thus $\phi({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})\leq\phi({\mathbf{T}}_{\star}^{(1)},{\mathbf{T}}_{\star}^{(2)})$ which implies that $({\mathbf{P}}^{(1)},{\mathbf{P}}^{(2)})$ is an optimal solution. Combining with eq. 27 we get
@@
where $f:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R},\ h:\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ and $g:\mathbb{R}^{N\times d}\to\mathbb{R}$ which is permutation invariant (Bronstein et al., 2021), then $\mathcal{C}$ is stable under permutation. Indeed, permuting the rows and columns of ${\mathbf{C}}_{Z}({\mathbf{Z}})$ by $\sigma$ is equivalent to considering the similarity ${\mathbf{C}}_{Z}({\mathbf{Y}})$, where ${\mathbf{Y}}=(\bm{z}_{\sigma(1)},\cdots,\bm{z}_{\sigma(n)})^{\top}$. Moreover, most similarities in the target space considered in DR take the form eq. 32: $\langle\Phi(\bm{z}_{i}),\Phi(\bm{z}_{j})\rangle_{\mathcal{H}}$ (kernels such as in spectral methods with $\Phi=\operatorname{id}$), $f(\bm{z}_{i},\bm{z}_{j})/\sum_{nm}f(\bm{z}_{n},\bm{z}_{m})$ (normalized similarities such as in SNE and t-SNE). Also note that the condition on $\mathcal{C}=\{{\mathbf{C}}_{Z}({\mathbf{Z}}):{\mathbf{Z}}\in\mathbb{R}^{N\times d}\}$ of proposition B.5 is met as soon as ${\mathbf{C}}_{Z}:\mathbb{R}^{N\times d}\to\mathbb{R}^{N\times N}$ is permutation equivariant.
Appendix C Generalized Semi-relaxed Gromov-Wasserstein is a divergence
Remark C.1 (Weak isomorphism).
According to the notion of weak isomorphism in Chowdhury and Mémoli (2019), for a graph $({\mathbf{C}},{\mathbf{h}})$ with corresponding discrete measure $\mu=\sum_{i}h_{i}\delta_{x_{i}}$, two nodes $x_{i}$ and $x_{j}$ are the “same” if they have the same internal perception i.e $C_{ii}=C_{jj}=C_{ij}=C_{ji}$ and external perception $\forall k\neq(i,j),C_{ik}=C_{jk},C_{ki}=C_{kj}$. So two graphs $({\mathbf{C}}_{1},{\mathbf{h}}_{1})$ and $({\mathbf{C}}_{2},{\mathbf{h}}_{2})$ are said to be weakly isomorphic, if there exist a canonical representation $({\mathbf{C}}_{c},{\mathbf{h}}_{c})$ such that $\operatorname{card}(\operatorname{supp}({\mathbf{h}}_{c}))=p\leq n,m$ and $M_{1}\in\left\{0,1\right\}^{n\times p}$ (resp. $M_{2}\in\left\{0,1\right\}^{m\times p}$) such that for $k\in\left\{1,2\right\}$
@@ Appendix C Generalized Semi-relaxed Gromov-Wasserstein is a divergence
We first emphasize a simple result extending a proof in Vincent-Cuaz et al. (2022a).
Proposition C.2.
@@
()($\Rightarrow$)( ⇒ ) As $\operatorname{srGW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$ there exists ${\mathbf{T}}\in\mathcal{U}({\mathbf{h}},\overline{{\mathbf{h}}})$ such that $E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=0$ hence $\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=0$. Using proposition C.2, $\operatorname{GW}_{L_{2}}=0$ hence using Theorem 18 in Chowdhury and Mémoli (2019), it implies that $({\mathbf{C}},{\mathbf{h}})$ and $(\overline{{\mathbf{C}}},\overline{{\mathbf{h}}})$ are weakly isomorphic.
@@
C.1 About trivial solutions of semi-relaxed GW when $L$ is not a proper divergence
@@
Appendix D Clustering properties: Proof of theorem 4.1
@@
We recall that $F({\mathbf{T}},\overline{{\mathbf{C}}}):=E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ and $G({\mathbf{T}})=F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))$. From lemma D.1 we know that $\overline{{\mathbf{C}}}({\mathbf{T}})$ is a minimizer of $\overline{{\mathbf{C}}}\to F({\mathbf{T}},\overline{{\mathbf{C}}})$ hence it satisfies the first order conditions $\partial_{2}F({\mathbf{T}},\overline{{\mathbf{C}}}({\mathbf{T}}))=0$. Every quantity is differentiable on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$. Hence, taking the derivative of $G$ and using the first order conditions
@@
We will found the expression of this gradient. In the proof of lemma D.1 we have seen that
@@
From lemma D.1 we have the expression (since ${\mathbf{P}}\in{\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$)
@@
Since ${\mathbf{U}}$ is symmetric we also have ${\mathbf{U}}\bm{1}_{N}={\mathbf{h}}_{X}$ and similarly we have the same result for ${\mathbf{V}}$. Overall ${\mathbf{V}}^{\top}\bm{1}_{N}={\mathbf{U}}^{\top}\bm{1}_{N}$ and ${\mathbf{V}}\bm{1}_{N}={\mathbf{U}}\bm{1}_{N}$. Since ${\mathbf{C}}_{X}({\mathbf{X}})$ is CPD or CND we can apply lemma D.3 below which proves that $-\operatorname{Tr}({\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))-\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}))+2\operatorname{Tr}({\mathbf{V}}{\mathbf{C}}_{X}({\mathbf{X}}){\mathbf{U}}{\mathbf{C}}_{X}({\mathbf{X}}))\leq 0$ and consequently that $G$ is concave on ${\mathcal{U}}^{+}_{n}({\mathbf{h}}_{X})$. We now use the continuity of $G$ to prove that it is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$.
@@
and hence $G$ is concave on ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$. This proves theorem 4.1. Indeed the minimization of ${\mathbf{T}}\in{\mathcal{U}}_{n}({\mathbf{h}}_{X})\to\min_{\overline{{\mathbf{C}}}}E_{L}({\mathbf{C}}_{X}({\mathbf{X}}),\overline{{\mathbf{C}}},{\mathbf{T}})$ is a minimization of a concave function over a polytope, hence admits an extremity of ${\mathcal{U}}_{n}({\mathbf{h}}_{X})$ as minimizer. But these extremities are membership matrices as they can be described as $\{\operatorname{diag}({\mathbf{h}}_{X}){\mathbf{P}}:{\mathbf{P}}\in\{0,1\}^{N\times n},{\mathbf{P}}^{\top}\bm{1}_{n}=\bm{1}_{N}\}$ (Cao et al., 2022). ∎
Lemma D.3.
@@
Appendix E Algorithmic details
@@
E.1 Objective function and gradient computation.
@@
Let consider any matrices ${\mathbf{C}}\in\mathbb{R}^{n\times n}$, $\overline{{\mathbf{C}}}\in\mathbb{R}^{m\times m}$, and a probability vector ${\mathbf{h}}\in\Sigma_{n}$. In all our use cases, we considered inner losses $L:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}_{+}$ which can be decomposed following Proposition 1 in Peyré et al. (2016). Namely we assume the existence of functions $f_{1},f_{2},h_{1},h_{2}$ such that
@@
$\begin{split}L_{2}(a,b)=(a-b)^{2}&\implies f_{1}(a)=a^{2},\quad f_{2}(b)=b^{2},\quad h_{1}(a)=a,\quad h_{2}(b)=2b,\\ L_{KL}(a,b)=a\log\frac{a}{b}-a+b&\implies f_{1}(a)=a\log a-a,\quad f_{2}(b)=b,\quad h_{1}(a)=a,\quad h_{2}(b)=\log b\\ L_{BCE}(a,b)=a\log\frac{a}{b}+(1-a)\log\frac{1-a}{1-b}&\implies f_{1}(a)=a\log a+(1-a)\log(1-a),\quad f_{2}(b)=-\log(1-b),\\ &\qquad h_{1}(a)=a,\quad h_{2}(b)=\log\frac{b}{1-b}\\ \end{split}$ (L2)
@@
Inspired from the work of Scetbon et al. (2022), we propose implementations of $\operatorname{srGW}$ that can leverage the low-rank nature of ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$. Let us assume that both structures can be exactly decomposed as follows, ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{A}}_{1}{\mathbf{A}}_{2}^{\top}$ where ${\mathbf{A}}_{1},{\mathbf{A}}_{2}\in\mathbb{R}^{N\times r}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})={\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top}$ with ${\mathbf{B}}_{1},{\mathbf{B}}_{2}\in\mathbb{R}^{n\times s}$, such that $r<<N$ and $s<<n$, can differ respectively from respective dimensions $p$ and $d$ (e.g. for used squared Euclidean distance matrices $r=p+2$ and $s=d+2$ ). For both inner losses $L$ we make use of the following factorization:
$L=L_{2}$: Computing the first term $F_{1}$ coming for the optimized second marginal can benefit from being factored if $d^{2}<<n$. Indeed, as $f_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))={\mathbf{C}}_{Z}({\mathbf{Z}})^{2}=({\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top})\odot({\mathbf{B}}_{1}{\mathbf{B}}_{2}^{\top})$, one can use the flattened out product operator described in Scetbon et al. (2022, Section 5), to compute ${\mathbf{C}}_{Z}({\mathbf{Z}})^{2}{\mathbf{T}}^{\top}\bm{1}_{N}={\mathbf{x}}$ in $O(min(n^{2},ns^{2}))$. This way $F_{1}({\mathbf{T}})$ results from stacking $N$ times ${\mathbf{x}}$ in $O(1)$ operations for a total number of computions of $N+O(min(n^{2},ns^{2}))$.And its scalar product with ${\mathbf{T}}$ to compute the loss comes down to $O(Nn)$ additional operations.
Then computing $F_{2}({\mathbf{T}})$ and its scalar product with ${\mathbf{T}}$ can be done following the development of Scetbon et al. (2022, Section 3) for the corresponding GW problem, in $O(Nn(r+s)+rs(N+n))$ operations. So the overall complexity at is $O(Nn(r+s)+rs(N+n)+min(n^{2},ns^{2}))$.
$L=L_{KL}$: In this setting $f_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))={\mathbf{C}}_{Z}({\mathbf{Z}})$ and $h_{1}({\mathbf{C}}_{X}({\mathbf{X}}))={\mathbf{C}}_{X}({\mathbf{X}})$ naturally preserves the low-rank nature of input matrices, but $h_{2}({\mathbf{C}}_{Z}({\mathbf{Z}}))=\log({\mathbf{C}}_{Z}({\mathbf{Z}}))$ does not. So computing the first term $F_{1}$, can be performed following this paranthesis order $\bm{1}_{N}((\bm{1}_{N}^{\top}{\mathbf{T}}){\mathbf{A}}_{1})){\mathbf{A}}_{2}^{\top})$ in $O(N(n+s))$ operations. While the second term $F_{2}$ should be computed following this order ${\mathbf{A}}_{1}(({\mathbf{A}}_{2}^{\top}{\mathbf{T}})\log({\mathbf{C}}_{Z}({\mathbf{Z}})))$ in $O(Nnr+rn^{2})$ operations. While their respective scalar product can be computed in $O(Nn)$. So the overall complexity is $O(Nnr+n^{2}r)$. Similar considerations can be applied to $L_{BCE}$.
@@
E.2 Solvers.
We develop next our extension of both the Mirror Descent and Conditional Gradient solvers first introduced in Vincent-Cuaz et al. (2022a), for any inner loss $L$ that decomposes as in equation 74 .
@@
where ${\mathbf{K}}^{(i)}=\exp\left(\ \nabla_{{\mathbf{T}}}F({\mathbf{T}}^{(i)})-\varepsilon\log({\mathbf{T}}^{(i)})\right)$ and $\varepsilon>0$ is an hyperparameter to tune. Proposition 3 and Lemma 7 in Vincent-Cuaz (2023, Chapter 6) provides that the Mirror-Descent algorithm converges to a stationary point non-asymptotically when $L=L_{2}$. A quick inspection of the proof suffices to see that this convergence holds for any losses $L$ satisfying equation 74, up to adaptation of constants involved in the Lemma.
@@
This algorithm, known to converge to local optimum (Lacoste-Julien, 2016), iterates over the 3 steps summarized in Algorithm 1: .
@@
Observe that this objective function can be developed as a second order polynom $g(\gamma)=a\gamma^{2}+b\gamma+c$. To find an optimal $\gamma$ it suffices to express coefficients $a$ and $b$ to conclude using Algorithm 2 in Vayer et al. (2018).
@@
Appendix F Appendix of experimental section